A slide at a park is 25ft long, an the top of the slide is 10ft above the ground. What is the approximate measure of the angle the slide makes with the ground?
Answer by stanbon(75887)

A slide at a park is 25ft long, an the top of the slide is 10ft above the ground. What is the approximate measure of the angle the slide makes with the ground?
----
Draw the picture.
height = 10
hypotenuse = 25
----
Let the angle be "x":
Equation:
sin(x) = 10/25 = 2/5
----
x = sin^-1(2/5) = 23.58 degrees = 23 degrees 35 minutes
==================
cheers,
Stan H.